population genetics

***Select the correct answer and explain why you have chosen this answer.

Sample Problem #4

In sheep, white vs yellow body fat is determined by a single pair of autosomal alleles. In a large, randomly mating flock, 23 percent of the sheep have white fat and 77 percent have yellow fat. You may conclude from this data that:

a. The allele for yellow is dominant to white

b. The allele for white is dominant to yellow

c. The alleles for yellow and white are codominant

d. The alleles for white and yellow are in a Hardy-Weinberg equilibrium

e. None of the above are valid conclusions
Answer #1

D) The alleles of white and yellow are in a Hardy-Weinberg Equilibrium (HWE).

It can be noticed here that the characters of heterozygotes (Wy and wY) are indiatinguisindis from dominant homozygote (WY), so they all are contributing to 77% of yellow fat colour.

Punnett Square
Y y
W WY Wy
w wY wy

Moreover, the frequency of allele for white is (23/100)=0.23

The frequency of allele yellow is (77/100)=0.77

Therefore, as per HWE: p+q=1 which implies; 0.23+0.77=1